# More Expressive Attention with Negative Weights

### v2

@@ 4 Experiments
We construct Transformer-like models using Cog Attention as the attention module, named Cogformer. We implement Cogformer in various tasks, including language modeling and image generation, to serve as a decoder-only model and a diffusion model, respectively. To showcase the effectiveness of Cog Attention  we evaluate and compare the performance of Cogformer with the Transformer across these diverse tasks.
@@ 4.1 Decoder-Only Models for Language Modeling
Architecture Details and Hyper-parameters. We train decoder-only Cogformer and Transformer-based language models on the RedPajama dataset (Computer, 2023). Apart from variations in the attention modules, the overall architecture and training hyper-parameters for Cogformer are identical to those of the vanilla Transformer. Importantly, we preserve the softmax attention in both the first and last layers of a deep Cogformer; we will elaborate on this design choice in Section 5.1. We use rotary position embedding (Su et al., 2023) and SwiGLU (Shazeer, 2020) activation in the feed-forward networks (FFN) following Llama (Touvron et al., 2023). RMSNorms are applied before both the attention and FFN modules. We adopt the Llama tokenizer, having a vocabulary of 32,000 tokens.
The model, with 141M parameters, consists of 12 layers, each with 12 attention heads. Its hidden state dimension is 768, while the intermediate dimension of the MLP layers is 3,072, with a context length of 2,048 tokens.
@@
The models are trained on 100 billion tokens, with batch size of 131,072 tokens, an initial learning rate of 2e-4, and linear warmup for the first 2,000 steps followed by a cosine decay schedule to 4% of the peak learning rate. The AdamW optimizer (Loshchilov & Hutter, 2019) is configured with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a norm clipping value of 1, and a weight decay of 0.1. Our training is conducted on 8$\times$×A800-80G GPUs for approximately one week.
Test Tasks. We evaluate language models on a variety of widely used tasks that comprehensively assess their language modeling capabilities. These tasks include:
(1) ARC-easy (Clark et al., 2018), a multiple-choice question-answering task, and ARC-challenge, which presents greater challenges than ARC-easy.
(2) PIQA (Bisk et al., 2020), a task on physical commonsense reasoning.
(3) SST-2 (Socher et al., 2013), a binary sentiment classification task.
(4) MNLI (Williams et al., 2018), where the model predicts whether a given pair of sentences entail, contradict, or are unrelated to each other.
(5) MRPC (Dolan & Brockett, 2005), where the model assesses whether two sentences have equivalent semantics.
(6) QQP  [^4] https://quoradata.quora.com/First-Quora-Dataset-Release-Question-Pairs, where the model determines whether two questions are semantically equivalent.
(7) RTE (GLUE (Wang et al., 2019) version), a two-class task for recognizing textual entailment.
All tasks are evaluated using accuracy as the metric, with all tests conducted in zero-shot, except for SST-2 which is one-shot tested. The evaluation code is based on LM Evaluation Harness (Gao et al., 2024).
Results. Table 1 presents the evaluation results. Cog Attention enhances the performance of Cogformer over the Transformer architecture across various tasks, highlighting its effectiveness in practical applications. Additionally, Figure 6 displays the training loss curves for the first 50,000 steps of both the Transformer and Cogformer architectures. The loss dynamics of the two models show no significant differences.
Figure 6: The convergence rates of Cogformer and the vanilla Transformer are nearly identical.
[…]
U-ViT-S/2 3.39 5.99
U-ViC-S/2 (Ours) 3.27 5.85
Table 2: Our U-ViC achieves better FID scores in unconditional image generation (CIFAR-10) and text-to-image generation (MS-COCO).
@@ 4.2 Diffusion Models for Image Generation
For our image generation tasks, we use U-ViT (Bao et al., 2023) as our baseline, which employs the Vision Transformer (Dosovitskiy et al., 2021) architecture as a diffusion model. We replace the softmax attention in U-ViT with Cog Attention (excluding the first and the last layers), resulting in our diffusion model, U-ViC. Both the U-ViC model and the U-ViT baseline are trained at a scale of 44 million parameters. The training is performed on the CIFAR-10 dataset (Krizhevsky, 2009) for unconditional image generation and the MS-COCO dataset (Lin et al., 2015) for text-conditioned image generation. For detailed information on the model architecture and training hyperparameters, please refer to Table 2 and Table 5 in (Bao et al., 2023). We evaluate performance using the FID score (Heusel et al., 2017), where lower values indicate better performance. Our experiments used 8 A800 GPUs, whereas Bao et al. (2023) used 4 × 2080 or 4 × A100 GPUs. Furthermore, we did not employ xFormer for acceleration. These differences in environments may contribute to the differences between the results reported in (Bao et al., 2023) and our reproduced ones.
Table 2 presents the results of our experiments. Our U-ViC model outperforms the U-ViT model across two datasets. Figure 7 illustrates generation samples.
Figure 7: Image generation samples. The figures are generated from the best checkpoint with lowest validation FID, and not cherry-picked.
@@ 5 Discussions
We discuss some important characteristics of Cog Attention, which might help the future research.
@@ 5.1 Convergence of Cogformer
As briefly mentioned before, in a deep Cogformer, we preserve the softmax attention in both the first and last layers. This aims to maintain the same convergence rate as a vanilla Transformer.
While we initially attempted to apply Cog Attention across all layers, we found that it resulted in slower convergence. Our observations of attention dynamics revealed that, at the beginning of training, the signs of attention weights do not accurately represent meaningful semantics due to the query-key inner product not being fully learned, leading to diverted optimization direction. The model appears to learn a shortcut at first, achieving lower training loss compared to the vanilla Transformer but then becoming stuck, allowing the vanilla Transformer to surpass it as training progresses. In contrast, softmax attention offers a near-uniform distribution that facilitates smoother initial training. Through empirical trials, we found that preserving softmax attention in the first layer of Cogformer effectively resolves the convergence issue.
Regarding the preservation of softmax attention in the last layer, we observed that a Cogformer with softmax attention only in the first layer resulted in last-layer attention weights’ sign highly consistent—either entirely positive or entirely negative. This suggests that the last layer may not require the expressiveness of negative weights. The training loss for these variants are shown in Table 3.
@@
$\text{Cogformer}_{\text{0}}$ 6.14 4.72 3.99 3.55 3.39
$\text{Cogformer}_{\text{0,last}}$ 6.15 4.73 4.02 3.53 3.38
Table 3: Training loss over the first 4500 steps for various models: $\text{Cogformer}_{\text{all}}$ (using Cog Attention in all layers), $\text{Cogformer}_{\text{0}}$ (applying softmax in the first layer), $\text{Cogformer}_{\text{0,last}}$ (applying softmax in both the first and last layers), and the vanilla Transformer (141M parameters).
@@ 5.2 Cog Attention Produces More Diverse Attention Patterns and Less Sink
Our analysis reveals that Cog Attention generates more diverse attention patterns compared to softmax attention. To investigate this, we examined the attention patterns across all heads in both Cogformer and Transformer, using the abstract section of “Attention Is All You Need” (Vaswani et al., 2017) as an input case. A comparison of Figures 9 and 10 shows that most heads in the vanilla Transformer exhibit sparse attention weights and a significant attention sink (Xiao et al., 2024), indicating that these heads are relatively inactive when processing the current input (Miller, 2023). In contrast, Cogformer displays more diverse attention patterns in its middle layers with reduced attention sink, suggesting that more heads are engaged in processing the input. We hypothesize that flexible attention patterns introduced by negative weights may lead to reduced parameter redundancy, but we currently lack direct evidence to quantify this reduction. Additionally, the diminished attention sink could potentially enhance extrapolation capabilities (Chen et al., 2024), KV cache compression (Liu et al., 2023b), high context-awareness task performance (Lin et al., 2024), and mitigate the lost-in-the-middle issue (Liu et al., 2023a). We will explore these questions in future research.
5.3 Time Cost Overhead
We compare the average time cost per training step on a single A800-80G GPU for our trained Transformer language model and Cogformer, using inputs of varying lengths. The results are shown in Figure 8. Due to the additional absolute operation and an extra multiplication, Cogformer incurs extra time overhead compared to the Transformer. We recognize this as a current limitation of our method and are actively exploring strategies to enhance its efficiency.
Figure 8: Comparison of time cost per training step for inputs of varying lengths between Transformer and Cogformer.
@@ 6 Related Works
Numerous efforts modified softmax function in the attention mechanism mainly for efficiency purposes, yet these attempts have not removed the constraints for non-negative weights. Some studies have proposed softmax alternatives such as ReLU (Shen et al., 2023; Wortsman et al., 2023), sigmoid (Ramapuram et al., 2024), cosine-based distance re-weighting (Qin et al., 2022), and learnable activations (Liu et al., 2024).
Recent studies showed the potential of an attention mechanism tailored to incorporate negative weights. Gong et al. (2024) demonstrate that removing the softmax function from the Mixture of Experts (MoE) router does not negatively impact performance. Ye et al. (2024) introduced differential attention, which calculates final attention weights by subtracting the outputs of two attention heads, thereby aiming to reduce noise. It might produce a small amount of negative attention weights as a by-product of the subtraction. It requires careful tuning of the scaling factor between the two heads and the inclusion of an additional normalization layer to ensure convergence. In this paper, we proposed Cog Attention, which introduces negative weights without requiring extra parameters or meticulous hyperparameter tuning. We demonstrated that Cog Attention increases the expressiveness of the attention mechanism.
@@ 7 Conclusions
In this paper, we introduced Cog Attention, a novel attention mechanism that incorporates negative weights. Our analysis, grounded in mechanistic interpretability, reveal that negative weights in attention computation enhance the expressiveness of neural networks and increase robustness against representational collapse. We subsequently implemented Cogformer, a new variant of the Transformer that integrates Cog Attention as its attention layer, to train language models and image generation diffusion models. The results indicate that Cog Attention can enhance Transformer performance across these tasks. We also discussed several properties of Cog Attention that could be beneficial for various applications, which we will explore in future works.
@@ Acknowledgements
We appreciate the valuable discussions with Shen Nie regarding the U-ViT models, and to Yining Qian, Jia-Nan Li, Songhao Wu, and Shuqi Li for their helpful writing suggestions.

### v3

@@ 4 Experiments
We construct Transformer-like models using Cog Attention as the attention module, named Cogformer. We implement Cogformer in various tasks, including language modeling and image generation, to serve as a decoder-only model and a diffusion model, respectively.
@@ 4.1 Decoder-Only Models for Language Modeling
Baselines. To show the effectiveness of Cog Attention beyond the vanilla Transformer, we compare Cogformer with several Transformer variants that might produce negative attention weights as a by-product of their design, though none are specifically tailored for this purpose.
(1) Differential attention (Ye et al., 2024) assumes attention calculations are noisy and aims to denoise them by subtracting outputs from two attention heads. This occasionally produces negative attention weights.
However, this method requires complex initialization and learnable coefficients to balance the two heads, along with additional normalization layers, for stability.
(2) Centered attention (Ali et al., 2023) attributes the representational collapse to the eigenvalues of the attention matrix being constrained within a limited interval. It mitigates this issue by shifting the row-wise attention sum from 1 to 0, which may lead to negative weights:
$\textbf{o}_{i}=\sum^{i}_{j=0}(\textbf{a}_{i,j}-\frac{1}{i})\cdot\textbf{v}_{j}.$
However, this shift depends on the input length rather than the input content, and does not fully remove the value constraints within the attention matrix, making the approach suboptimal. Notably, this work can provide additional theoretical insights into the benefits of Cog Attention.
General Implementation and Hyper-parameters. We train decoder-only Transformer language models and their variants using Cog Attention and its baselines on the RedPajama dataset (Computer, 2023). Apart from differences in the attention modules, the overall architecture and training hyperparameters remain consistent across all models.
Following the Llama (Touvron et al., 2023), we employ rotary position embedding (Su et al., 2023) and SwiGLU activation (Shazeer, 2020) in the feed-forward networks (FFN). RMSNorm (Zhang & Sennrich, 2019) is applied prior to both the attention and FFN modules. The Llama tokenizer, with a vocabulary of 32,000 tokens, is used. The models have 141 million parameters, comprising 12 layers, each with 12 attention heads. The hidden state dimension is 768, and the intermediate dimension of the MLP layers is 3,072. The training context length is 2,048 tokens.
@@
We use a batch size of 64, an initial learning rate of 2e-4, and linear warm-up for the first 2,000 steps followed by a cosine decay schedule to 4% of the peak learning rate (Tow et al., 2024). The AdamW optimizer (Loshchilov & Hutter, 2019) is configured with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a norm clipping value of 1, and a weight decay of 0.1. Each model is trained on 100 billion tokens, using 8$\times$×A800-80G GPUs for approximately one week.
Model-Specific Settings. Differential attention requires a complex initialization of relative scales between attention heads. We follow the configuration described in the original paper; for details, please refer to (Ye et al., 2024). For Cogformer, we preserve softmax attention in both the first and last layers. This is design choice is explained in Section 5.1. Centered attention does not need special settings.
Table 1: Performance comparison across multiple NLP tasks among language models using different attention algorithms. We highlight the top and second-best results in each task with bold text and underlining, respectively. Cogformer achieves the highest average accuracy compared to other Transformer variants, which may occasionally produce negative attention as a by-product of specific operations.
Model ARC-E ARC-C PIQA SIQA MRPC SST2 MNLI Avg.
$\text{Transformer}_{\text{141M}}$ 42.34 19.54 57.73 37.00 60.54 51.72 32.05 42.98
w/ Differential Attn. 40.78 18.86 57.89 37.10 58.82 53.56 34.38 43.06
w/ Centered Attn. 40.66 19.11 58.16 36.39 58.33 55.28 32.76 42.96
$\text{Cogformer}_{\text{141M}}$ (Ours) 43.90 19.54 59.09 37.36 62.25 54.59 33.71 44.35
Table 2: Larger Cogformer models (600M parameters on 250B tokens and 1B parameters on 50B tokens) outperform Transformer models.
Model ARC-E ARC-C PIQA SIQA MRPC SST2 MNLI Avg.
$\text{Transformer}_{\text{600M}}$ 51.01 21.33 64.15 40.28 58.09 54.09 33.49 46.06
$\text{Cogformer}_{\text{600M}}$ 52.40 23.46 63.71 40.51 67.16 55.05 33.56 47.97
$\text{Transformer}_{\text{1B}}$ 47.81 20.22 61.75 37.22 58.33 34.81 58.49 45.52
$\text{Cogformer}_{\text{1B}}$ 48.86 21.59 62.24 38.54 58.33 32.24 61.70 46.21
Evaluation. We evaluate language models across a range of widely used tasks. Specifically, we assess four reasoning tasks: (1) grade-school science, using ARC-easy and ARC-challenge (Clark et al., 2018); (2) physical commonsense, using PIQA (Bisk et al., 2020); (3) social situations, using SIQA (Sap et al., 2019). Moreover, following (Wang et al., 2019), we evaluate language models (one-shot) on single-sentence tasks, similarity and paraphrase tasks, as well as natural language inference (NLI) tasks. These include: (4) SST-2 (Socher et al., 2013) for binary sentiment classification, (5) MRPC (Dolan & Brockett, 2005) for assessing semantic equivalence between sentence pairs, and (6) MNLI (Williams et al., 2018) for determining entailment, contradiction, or neutrality between sentence pairs. The evaluation code is based on the LM Evaluation Harness (Gao et al., 2024), and the metric used is accuracy.
Results. Table 1 presents the evaluation results, showing that Cogformer achieves stable improvements over the vanilla Transformer in most tasks, with the highest overall average accuracy. While our baseline models achieve top results in several tasks, their performance is unstable, with some tasks experiencing drops in accuracy compared to the vanilla Transformer. Meanwhile, the fluctuating impact on performance observed with Differential Attention further indicates that the additional coefficients introduced to stabilize training may not be robust enough across different training datasets or model architectures. In contrast, Cog Attention does not introduce any additional modules or (un)learnable coefficients, ensuring effective training without added complexity.
Scalability. To demonstrate the scalability of Cog Attention, we train larger models with 600 million parameters on 250 billion tokens, and 1 billion parameters on 50 billion tokens. Detailed architecture and training hyperparameters are provided in Appendix C. Due to the prohibitive cost of implementing all baselines at this scale, we only compare Cogformer with the vanilla Transformer. Meanwhile, training models at a larger scale or using more tokens is unfeasible for us. As shown in Table 2, larger Cogformers still outperforms Transformers. Figure 8 in the appendix presents performance across tasks for 600M models trained with 100, 150, and 200 billion tokens, where Cogformer achieves better results in most comparisons. These results highlight the potential of Cog Attention for pretraining advanced large language models.
@@ 4.2 Diffusion Models for Image Generation
We adopt U-ViT (Bao et al., 2023) as the backbone model, which leverages the Vision Transformer (Dosovitskiy et al., 2021) architecture as a diffusion model. We enhance U-ViT by replacing its softmax attention (except in the first and last layers) with Cog Attention, resulting in our modified diffusion model, U-ViC. The remaining baselines from the previous section are not compared, as they were originally designed and evaluated only on NLP tasks.
All models have 44 million parameters. Training is performed on the CIFAR-10 dataset (Krizhevsky, 2009) for unconditional image generation and the MS-COCO dataset (Lin et al., 2015) for text-conditioned image generation. For detailed information on the model architecture and training hyperparameters, please refer to Table 2 and Table 5 in (Bao et al., 2023). We evaluate performance using the FID score (Heusel et al., 2017), where lower values indicate better performance. All models are trained on 8$\times$×A800 GPUs. Some environment differences may lead to slightly different reproduced results. Table 3 presents the results. Our U-ViC model shows superior performance across both tasks. Figure 9 illustrates generated samples.
Table 3: Our U-ViC achieves lower FID in unconditional image generation (CIFAR-10) and text-to-image generation (MS-COCO).
Model CIFAR-10 MS-COCO
U-ViT 3.39 5.99
U-ViC (Ours) 3.27 5.85
@@ 5 Discussions
We discuss some important characteristics of Cog Attention, which may help future research.
@@ 5.1 Convergence of Cogformer
As briefly mentioned earlier, in a deep Cogformer, we preserve the softmax attention in both the first and last layers. This aims to maintain, or even improve, the convergence rate compared to a vanilla Transformer.
We initially attempted to apply Cog Attention across all layers, but this resulted in slower convergence. Our observations of attention dynamics revealed that, at the beginning of training, the signs of the attention weights do not accurately represent meaningful semantics due to the query-key inner product not being fully learned, which leads to diverted optimization direction. The model appears to learn a shortcut at first, achieving lower training loss compared to the vanilla Transformer but then becomes stuck, allowing the vanilla Transformer to surpass it as training progresses. In contrast, softmax attention offers a near-uniform distribution that facilitates smoother initial training. Through empirical trials, we found that preserving softmax attention in the first layer of Cogformer largely resolves the convergence issue. Regarding the preservation of softmax attention in the last layer, we observed that the signs of the attention weights in the last layer are highly consistent—either completely positive or completely negative. This suggests that the last layer may not require the expressiveness of negative weights. The training losses for these variants are presented in Table 4 where $\text{Cogformer}_{\text{all}}$ begins to exhibit higher losses after 2,500 steps.
Figure 6 shows the training losses for the models in our main experiments. When softmax attention is applied in the first and last few layers, Cogformer demonstrates more effective training across different scales, achieving a generally lower loss compared to the Transformer.
Table 4: Training loss over the first 4,500 steps for various models: $\text{Cogformer}_{\text{all}}$ (using Cog Attention in all layers), $\text{Cogformer}_{\text{0}}$ (applying softmax in layer 0), $\text{Cogformer}_{\text{0,last}}$ (applying softmax in both the first and last layers), and the vanilla Transformer (141M).
@@
$\text{Cogformer}_{\text{0}}$ 6.14 4.72 4.02 3.55 3.39
$\text{Cogformer}_{\text{0,last}}$ 6.15 4.73 3.99 3.53 3.38
Figure 6: Training losses for models at different scales over the first 100 billion tokens (50 billion for the 1B-parameter models).
@@ 5.2 Cog Attention Produces More Diverse Attention Patterns and Less Sink
Our analysis reveals that Cog Attention generates more diverse attention patterns compared with softmax attention. To investigate this, we examined the attention patterns across all heads in both $\text{Cogformer}_{\text{141M}}$ and $\text{Transformer}_{\text{141M}}$, using the abstract section of “Attention Is All You Need” (Vaswani et al., 2017) as an input case. A comparison of Figures 10 and 11 shows that most heads in the vanilla Transformer exhibit sparse attention weights and a significant attention sink (Xiao et al., 2024), indicating that these heads are relatively inactive when processing the current input (Miller, 2023). In contrast, Cogformer displays more diverse attention patterns in its middle layers with a reduced attention sink, which suggests that more heads are engaged in processing the input. We hypothesize that flexible attention patterns introduced by negative weights may lead to reduced parameter redundancy, but we currently lack direct evidence to quantify this reduction. Additionally, the diminished attention sink could potentially enhance extrapolation capabilities (Chen et al., 2024), KV cache compression (Liu et al., 2023b), high-context-awareness task performance (Lin et al., 2024), and mitigate the lost-in-the-middle issue (Liu et al., 2023a). We will explore these questions in future research.
@@ 6 Related Works
Numerous efforts have modified the softmax function in the attention mechanism mainly for efficiency purposes, yet these attempts have not removed the constraints on non-negative weights. Some studies have proposed softmax alternatives such as ReLU (Shen et al., 2023; Wortsman et al., 2023), sigmoid (Ramapuram et al., 2024), cosine-based distance re-weighting (Qin et al., 2022), and learnable activations (Liu et al., 2024). The approaches discussed in the “Baselines” paragraph of Section 4, which occasionally produce negative weights as a by-product of their specialized operations, showed the potential of an attention mechanism tailored to incorporate negative weights. In this paper, we propose Cog Attention, which introduces negative weights for increased expressiveness without requiring additional parameters or meticulous hyperparameter tuning—unlike the approaches discussed in Section 4. A Cogformer requires the maintenance of softmax attention in both the first and last few layers, highlighting the distinct characteristics of these layers, whose unique behavior has been discussed in several studies (Gong et al., 2024; Cancedda, 2024).
@@ 7 Conclusions
We introduce Cog Attention, a novel attention mechanism that incorporates negative weights. Mechanistic interpretations reveal that negative attention weights enhance the expressiveness of Transformers and increase robustness against representational collapse. We train Cogformer, a new variant of the Transformer that integrates Cog Attention as its attention layer, as both language models and image generation diffusion models. Cogformer outperforms Transformers on these tasks. We also discuss several properties of Cog Attention that could be beneficial for various applications, which we will explore in future work.
@@ Acknowledgements
We appreciate the valuable discussions with Shen Nie about the U-ViT models and to Yining Qian, Jia-Nan Li, Songhao Wu, and Shuqi Li for their helpful writing suggestions.
Appendix A Efficiency Analysis
We compare the efficiency of Cog Attention and softmax attention, both implemented using Triton fused kernels, on a single A800-80G GPU. Figure 7 illustrates the throughput and memory costs for Cog Attention and softmax attention, given a QK-product matrix of size $T\times T$. Compared to softmax attention, Cog Attention introduces no additional time overhead and a negligible increase in memory usage when $T$ is very large.
Figure 7: Cog Attention is as efficient as softmax attention.
Appendix B How Do Negative Attention Weights Mitigate the “Over-Squashing” Issue
Table 5: An overview of symbols used in the theorem for convenience.
Symbol Explanation
$L$ Maximum number of layers in the model.
$l$ Index of a layer, ranging from 00 to $L$; layer 00 represents the embedding layer.
$\textbf{v}^{l}_{i}$ The $i$-th input at the $l$-th layer.
$\textbf{y}_{i}$ The $i$-th output of the entire model.
$\alpha^{l}_{j,i}$ Attention weight from position $j$ to position $i$ at layer $l$.
$\sigma$ The maximum Lipschitz constant of any MLP layer.
$\sigma^{l}$ Lipschitz constant of the MLP at layer $l$.
$\beta^{l}_{i}$ The effect of $i$-th normalization module at layer $l$, where $i\in\{1,2\}$
$\beta_{3}$ The effect of normalization module applied after all layers.
$k_{l}$ Enumeration index along the sequence length dimension within layer $l$.
We explain Cogformer’s enhanced robustness against representational collapse using the theorem B.5 from (Barbero et al., 2024). The theorem states that for an input sequence $\mathbf{v}^{0}_{1},\dots,\mathbf{v}^{0}_{n}$, the following inequality holds:
$\lVert\frac{\partial\mathbf{y}_{n}}{\partial\mathbf{v}^{0}_{i}}\rVert\leq C\sum_{k_{1}\geq i}\dots\sum_{k_{L}\geq k_{L-1}}\bar{\alpha}^{L-1}_{n,k_{L}}\prod_{l=2}^{L-1}\bar{\alpha}^{l-1}_{k_{l},k_{l-1}}\bar{\alpha}^{0}_{k_{1},i},$
where $\bar{\alpha}^{l}_{j,i}=\lvert\frac{\alpha^{l}_{j,i}}{\beta^{l}_{1}}+\mathbbm{1}(i=j)\rvert$ and $C=\frac{1}{\beta_{3}}\prod_{l=1}^{L}\left(\frac{\sigma}{\beta^{l}_{2}}+1\right)$.
We can see that allowing attention weights to be negative reduces the upper bound of $\left\|\frac{\partial\mathbf{y}_{n}}{\partial\mathbf{v}^{0}_{i}}\right\|$. Specifically, for any layer $l$, when $k_{l}=k_{l-1}$, the expectation of $\bar{\alpha}^{l}_{k_{l},k_{l-1}}$ decreases, thereby lowering the right side of the inequality. An intuitive explanation is that negative weights can adaptively diminish the information flow from position $k_{l-1}$ in layer $l-1$ to the same position $k_{l-1}$ in layer $l$. Consequently, for any $k_{l+1}\geq k_{l-1}$, $\mathbf{v}^{l}_{k_{l+1}}$ is less likely to be over-squashed by $\mathbf{v}^{l}_{k_{l-1}}$. By “adaptively diminish,” we mean that when $\mathbf{v}^{0}_{i}$ contains irrelevant information (e.g., token $i$ is “1” while the query is “To find a zero”), Cogformer’s attention to token $i$ tends to be negative, mitigating the over-squashing of irrelevant information. This enables the model to focus on useful tokens.
It is important to note that the information in $\mathbf{v}^{0}_{i}$ is not only semantic but also positional. For tasks where tokens are semantically similar (e.g., task 2 in Section 3.2), the position information from earlier tokens is less over-squashed into the outputs, allowing more information pathways for diverse positions. For reference, several works have studied the critical role of positional information in Transformers’ counting ability (Barbero et al., 2024; Yehudai et al., 2024; Golkar et al., 2024).
Appendix C Implementation Specifics and Hyper-parameters for Larger Models
The architectures of our middle and large models share the same position embedding, FFN structure, and vocabulary as the small language models (141M). The training context length is set to 4,096 tokens.
The training hyper-parameters follow those outlined in (Tow et al., 2024): a batch size of 4 million tokens, an initial learning rate of $3.2\times 10^{-4}$, and a linear warmup over the first 4,800 steps, followed by a cosine decay schedule that reduces the learning rate to 4% of its peak value. The AdamW optimizer (Loshchilov & Hutter, 2019) is used with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a gradient norm clipping threshold of 1, and a weight decay of 0.1.
A 600M-parameter model consists of 16 layers, with 24 attention heads per layer. The hidden state dimension is 1,536, and the intermediate dimension of the MLP layers is 4,608. The first and last two layers use softmax attention. Training a model (with 250B tokens) at this scale requires approximately 3,600 GPU hours (A800-80G).
A 1B-parameter model consists of 22 layers, with 28 attention heads per layer. The hidden state dimension is 1,792, and the intermediate dimension of the MLP layers is 5,376. The first and last three layers use softmax attention. Training a model (with 50B tokens) at this scale requires approximately 1,600 GPU hours (A800-80G).
Figure 8: Performance across tasks for models trained with 100, 150, and 200 billion tokens.
Figure 9: Text-to-image generation samples on MS-COCO. The figures are generated by the best checkpoint with lowest validation FID, and not cherry-picked.
Figure 10: Attention patterns obtained from Transformer.
Figure 11: Attention patterns obtained from Cogformer.